OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 4y2

OpenStudy (mertsj):

How does this open...right, left, up, or down?

OpenStudy (anonymous):

up i think

OpenStudy (mertsj):

What evidence do you have for that?

OpenStudy (anonymous):

put the equation into the graphing caculator and it shows that the parabola opens upward

OpenStudy (mertsj):

Something is wrong then. Solve it for y and then put it into your calculator. Because this opens to the right.

OpenStudy (mertsj):

y=x^2 opens up x=y^2 opens to the right

OpenStudy (mertsj):